# You Only Look at Once for Real-time and Generic Multi-Task

### v1

@@ Abstract
High precision, lightweight, and real-time responsiveness are three essential requirements for implementing autonomous driving. Considering all of them simultaneously is a challenge. In this study, we present an adaptive, real-time, and lightweight multi-task model designed to concurrently handle object detection, drivable area segmentation, and lane detection tasks. To achieve this research objective, we developed an end-to-end multi-task model with a unified and streamlined segmentation structure. Our model operates without the need for any specific customization structure or loss function. We achieved competitive results on the BDD100k dataset, particularly in visualization outcomes. The performance results show a mAP50 of 81.1% for object detection, a mIoU of 91.0% for drivable area segmentation, and an IoU of 28.8% for lane line segmentation. Additionally, we introduced a real-road dataset to evaluate our model’s performance in a real scene, which significantly outperforms competitors. This demonstrates that our model not only exhibits competitive performance but is also more flexible and faster than existing multi-task models. The source codes and pre-trained models are released at https://github.com/JiayuanWang-JW/YOLOv8-multi-task
@@ I Introduction
Owing to rapid advancements in deep learning, the field of computer vision has flourished, particularly in autonomous driving applications. Autonomous driving systems (ADS) provide enhanced convenience for everyday driving. It effectively alleviates driver fatigue and reduces traffic accidents. Gaining traction in both consumer vehicles and commercial trucks, ADS aim to ensure a driving experience that is safe, comfortable, and efficient [1]. ADSs can either rely solely on multiple cameras and computer vision technology, or integrate an array of advanced sensors including lidars, cameras, and radars. [2]. Both approaches have their advantages. Compared with cameras, lidar sensors are less sensitive to light and night effects. Besides, it can provide 3D information with a large view [3]. However, the long development cycle and high costs of lidar sensors have hindered their widespread adoption. Conversely, using cameras has the advantage of being cost-effective and easy to install. When considering solely camera-captured images as input and combining them with deep learning models into ADSs, it becomes imperative to meet the stringent requirements of real-time processing. Specifically, maintaining a frame rate exceeding 30 frames per second (FPS) becomes imperative [4]. Quicker responses increase the likelihood of avoiding accidents and ensuring survival. Given the limited computational capacity of edge devices, achieving real-time performance and high precision with deep learning in autonomous driving is challenging.
@@
Numerous methods have been proposed for individual tasks in autonomous driving, many of which have achieved outstanding results. In the realm of object detection tasks, two primary methodologies are prevalent. The first category encompasses two-stage techniques, epitomized by the seminal Fast R-CNN [5]. While these methods prioritize detection accuracy, they often come at the expense of computational efficiency. The second category is represented by one-stage approaches, spearheaded by the You Only Look Once (YOLO) series [6, 7, 8, 9], ranging from version 1 to the latest, version 8. YOLOv8 is notable for its real-time detection capabilities and has been widely adopted for various detection tasks. It primarily focuses on detection rather than segmentation. While they introduced a segment head for segmentation tasks, the loss functions and evaluation metrics employed are toward the detection task. Additionally, yolov8 can only implement one task per model. In this case, YOLOv8 is not suitable for autonomous driving tasks. Implementing it under multi-task conditions requires multiple models, which significantly increases both training and inference times. In the context of autonomous driving, one crucial metric is inference time or FPS.
@@
In the context of the segmentation task, one of the field’s milestones is Fully Convolutional Networks (FCN) [10]. Additionally, U-Net [11] and SegNet [12] are common models in drivable area segmentation tasks. However, lane line segmentation differs from drivable area segmentation because of the distinct elongated and narrow features of lane lines in road images. Lane line segmentation necessitates low-level and multiple-scale features for effective analysis. Recently, models like PointLaneNet [13] and MFIALane [14] have gained popularity in the field of lane line segmentation. Although they have achieved outstanding results in their respective domains, integrating them presents challenges, primarily due to the differing resolution of features required. Segmentation operates at the pixel level, whereas detection tasks employ grid cells in one-stage methods and utilize selective search in two-stage approaches. While their focus differs, both segmentation and detection tasks necessitate initial feature extraction from input images. Consequently, sharing a backbone is possible. Compared to using separate models for each task, integrating three distinct necks and heads into a model with a shared backbone can save lots of computing resources and inference time.
@@
Recently, several multi-task models have been proposed and applied in the ADS, such as YOLOP [15], Shokhrukh [16] et al. proposed multi-task learning model, Sparse U-PDP [17] and HybridNet [18]. All of these methods selected three tasks to construct the panoptic autonomous driving system: object detection, drivable area segmentation, and lane line segmentation, which are from the public Berkeley Deep Drive(BDD100K) dataset. Each of these methods has yielded outstanding results. Typically, their models are composed of two components: an encoder and a decoder. However, several challenges remain to be addressed. First, they have complex structures in their neck or head components, which can further impact inference times. Second, they often focus on specialized tasks when designing the loss function or neck structure. As a result, the model might compromise its generality. Thus, developing a fast, robust, and universally applicable model is crucial.
Combining three heads into one model is a significant challenge because they rely on a shared backbone, and this backbone must be able to capture features for all three tasks simultaneously. In this case, the experiments may excel in one or several tasks, but some tasks may produce notably poor results. One possible solution is to tailor a specific loss function for each task. However, this approach lacks generalization. When we introduce a new task, we must modify the model accordingly, such as YOLOP [15], Shokhrukh [16] et al. proposed specifically loss functions or neck structure individually for drivable area and lane line segmentation. It is not friendly for practical applications. Therefore, developing a generic model remains a challenge.
@@
In our research, we proposed A-YOLOM, an adaptive model specifically designed for multi-task. Notably, A-YOLOM can efficiently handle multiple tasks with one model. Furthermore, by employing consistent loss functions across the same task types, we maintain a uniform approach. Importantly, we introduced an adaptive module tailored for the segmentation neck, thus eliminating the need for distinct designs for different scene tasks. To summarize, the primary contributions of our study are as follows:
@@
  • We have developed a lightweight model capable of integrating three tasks into a single unified model. This is particularly beneficial for multi-task that demand real-time processing.
@@
  • We have designed a novel Adaptive Concatenate Module specifically for the neck region of segmentation architectures. This module can adaptively concatenate features, further enhancing the model’s generality.
@@
  • We designed a lightweight, and generic segmentation head. No trick, simple and useful. We have a unified loss function for the same type of task head, meaning we don’t need to custom design for specific tasks.
@@
  • We evaluated our model on publicly available autonomous driving datasets, achieving competitive results, particularly in terms of inference time and visualization. Moreover, we further assessed our model’s performance using real road datasets and significantly outperformed competitors.
@@ II Related Works
In this section, we review past work on detection, segmentation, and multi-task models in autonomous driving tasks. We focus on deep learning-based methods.
@@ II-A Detection
Over the past decade, the swift advancements in computer vision have significantly bolstered the progression of autonomous driving. Specifically, these tasks can be delineated into two primary parts: detection and segmentation. The detection task encompasses object detection, identifying entities such as vehicles, persons, traffic signs, etc. The current object detection methods have been divided into two classifications [19]: two-stage and one-stage.
Two-stage methods initiate with a Region Proposal Network (RPN) to generate Regions of Interest (RoI). Subsequently, the second phase employs a deep learning network to classify these RoIs. This latter stage also fine-tunes the bounding box dimensions and position, bolstering object localization accuracy. In the two-stage object detection paradigm, notable methods include Regions with CNN features (R-CNN) [20], Fast R-CNN [5], and Faster R-CNN [21].
In contrast, the one-stage detection approach offers an end-to-end strategy. It simultaneously predicts the bounding box and classifies the object in a single forward pass. It can be deployed on mobile devices for real-time operation and is easily implementable. The notable methods include You Only Look Once (YOLO) [6, 7, 8, 9] series, Single-Shot Multibox Detector (SSD) [22], and RetinaNet [23].
While one-stage methods typically lag behind two-stage methods in terms of performance, their real-time capabilities have made them increasingly popular in object detection tasks.
@@ II-B Segmentation
The segmentation task, including semantic and instance segmentation, differs from detection as it typically operates at the pixel level. Common segmentation tasks in autonomous driving include drivable area detection and lane line segmentation. However, these are distinct tasks. The drivable area typically covers a large portion of an image, but numbers infrequently. Lane lines are the opposite: they are numerous in one image, yet each line is small, elongated, and narrow.
In 2015, the deep learning model Fully Convolutional Networks (FCN) [10] was introduced for semantic segmentation tasks. With its end-to-end training capability, it achieved a 20% improvement on the Pascal VOC 2012 dataset compared with traditional methods. Due to its focus on local information, global details are lost, which can produce coarse segmentation results. Therefore, the encoder-decoder architecture model was proposed. Badrinarayanan et al. [12] proposed SegNet, which preserves the maximum pooling indices to ensure more accurate detail restoration during the upsampling process. However, the encoder-decoder architecture has a limitation. High-resolution details are often lost during the encoding process, reducing fine-grained information. This is not ideal for tasks such as lane line segmentation. Pan et al. [24] proposed SCNN, which employs spatial convolutions to capture the continuity and structural information in images across both horizontal and vertical directions. Due to its ability to capture spatial correlations in images, it is well-suited for detecting slender objects, like lane lines. However, such convolution operations are computationally intensive and time-consuming.
Given the constraints of edge devices in vehicles, it’s crucial to consolidate the multi-task into a single model. This not only conserves computational resources but also meets real-time performance requirements.
@@ II-C Multi-task model
Recently, multi-task models have become a research focal point due to their high efficiency. Given that autonomous driving typically involves various sub-tasks and the compute resource limitation of edge devices, the multi-task model advantage becomes particularly pertinent. MultiNet [25] introduced an efficient, unified deep architecture capable of simultaneously addressing classification, detection, and semantic segmentation. YOLOP [15] extends the YOLOv5 model by introducing two additional segmentation necks and heads to achieve both segmentation and detection tasks simultaneously. They separately design the structure and loss function for each segmentation task to enhance performance, which will compromise the generality of the model. Sparse U-PDP [17] proposed a unified decoding framework that integrates three tasks into a single model. Notably, they crafted a streamlined multi-task representation through ”dynamic convolution kernels”, enhanced by a dynamic interaction module which adjusts feature sampling uniquely for each task. Due to its complex decoder, it cannot meet real-time requirements. Shokhrukh et al. [16] introduced a memory-efficient end-to-end framework that utilizes combinations of customized loss functions, encompassing the weighted average sum of all tasks to enhance performance. The customization and specific weighting will limit the model’s adaptability to new tasks. Additionally, their approach involves a sophisticated training paradigm. We critique complex algorithms because they are prone to overfit and maybe display subpar performance on real roads.
Therefore, developing a generic and real-time multi-task model is still a significant challenge. In our work, we introduce an adaptive concatenation module to enhance the model’s generality, and employ a lightweight backbone, and a simple and useful segmentation head to decrease parameter count, ultimately fulfilling real-time requirements.
@@ III PROPOSED METHODOLOGY
In this section, we will present the details of our model. Specifically, our discussion will encompass three main components: the backbone, neck, and head. Additionally, the loss function is also included. A-YOLOM structure is shown in Fig 1. Our model is a one-stage network with a simple encoder-decoder architecture. The encoder comprises both the backbone and neck, while the decoder consists of the head. It’s worth noting that we have three necks in total: one detection neck for object detection and two segmentation necks, one for drivable areas and another for lane lines. Our model boasts a simple and efficient structure, which not only broadens its application range but also ensures real-time inference performance.
@@ III-A Encoder
In our work, we utilize a shared backbone and employ three neck networks for the three different tasks.
@@ III-A1 Backbone
The backbone comprises a series of convolutional layers dedicated to extracting features from the input data. Given the outstanding performance of YOLOv8 in detection tasks, our backbone network follows YOLOv8. Specifically, they refined CSP-Darknet53 [26], which served as the backbone of YOLOv5. The primary differences in this backbone compared to CSP-Darknet53, are the use of c2f layers in place of the c3 module. The c2f module combines the high-level features with contextual information, further improving performance. There are also minor differences, such as using a 3x3 convolutionalin place of the initial 6x6 convolutional, and the removal of the 10th and 14th convolutional layers. These enhancements render the backbone more effective than its predecessors in the YOLO series.
@@ III-A2 Neck
The neck is responsible for fusing the features extracted from the backbone. After the backbone network, we use an SPPF module to increase the receptive field and reduce computational requirements compared to SPP. These features are subsequently directed to individual necks.
In our model, we utilize three necks: one for the object detection task and two for segmentation tasks, specifically for drivable areas and lane lines. It’s worth noting that in Fig 1, due to we have two segmentation tasks, the value of N in the segment neck on the left part is equal to 2. This means we have two same segment necks for different segmentation tasks. The sky-blue and brown lines, which originate from different neck bottom layers, lead to separate segment heads.
For the detection neck, we adopt the path aggregation network (PAN) [27] structure, which includes top-down and bottom-up two FPNs. This structure merges low-level details with high-level semantic features, enriching the overall feature representation. This is invaluable for object detection, where small objects hinge on low-level features, while larger entities gain from the broader context provided by high-level features. In detection tasks, objects of varying scales are commonly present. This diversity in object size is the primary reason we opted for the PAN structure for the detection neck.
For the segmentation neck, we adopt the Feature Pyramid Network (FPN) [28] structure, renowned for its effective handling of multi-scale objects. Additionally, we introduced an adaptive concatenation module between the neck and backbone. As illustrated in Algorithm 1, this module operates at each identical resolution level, adaptively determining whether features should be concatenated. This addition enhances the model’s generality, making it suitable for various segmentation tasks.
@@
0:  List of tensors $x$ includes two feature maps, channels list $ch$

### v4

@@ Abstract
High precision, lightweight, and real-time responsiveness are three essential requirements for implementing autonomous driving. In this study, we incorporate A-YOLOM, an adaptive, real-time, and lightweight multi-task model designed to concurrently address object detection, drivable area segmentation, and lane line segmentation tasks. Specifically, we develop an end-to-end multi-task model with a unified and streamlined segmentation structure. We introduce a learnable parameter that adaptively concatenates features between necks and backbone in segmentation tasks, using the same loss function for all segmentation tasks. This eliminates the need for customizations and enhances the model’s generalization capabilities. We also introduce a segmentation head composed only of a series of convolutional layers, which reduces the number of parameters and inference time. We achieve competitive results on the BDD100k dataset, particularly in visualization outcomes. The performance results show a mAP50 of 81.1% for object detection, a mIoU of 91.0% for drivable area segmentation, and an IoU of 28.8% for lane line segmentation. Additionally, we introduce real-world scenarios to evaluate our model’s performance in a real scene, which significantly outperforms competitors. This demonstrates that our model not only exhibits competitive performance but is also more flexible and faster than existing multi-task models. The source codes and pre-trained models are released at https://github.com/JiayuanWang-JW/YOLOv8-multi-task
@@ I Introduction
Owing to rapid advancements in deep learning, the field of computer vision has flourished, particularly in autonomous driving applications [1, 2, 3]. Autonomous driving systems (ADS) provide enhanced convenience for everyday driving. Detection and segmentation are crucial for ADS, which encompasses three core tasks: object detection, drivable area segmentation, and lane line segmentation, as shown in Figure 1. In autonomous driving tasks, lidar and camera sensors are often used to acquire environmental information. However, the camera-based method stands out due to its low cost. Therefore, combining cameras with a deep learning model becomes a formidable solution. Additionally, high precision, real-time and lightweight are essential in autonomous driving tasks. In emergencies, ADS must fast and accurately decide to avoid potential collisions or safely around obstacles. Accurately and fast estimating the drivable area and lane lines is important for effective route planning. In fact, maintaining a frame rate exceeding 30 frames per second (FPS) is imperative for ADS [4, 5]. Given the limited computational capacity of edge devices, employing lightweight models is necessary. However, achieving both real-time performance and high precision simultaneously in autonomous driving through a lightweight model poses significant challenges.
Figure 1: Multi-task in autonomous driving: object detection, drivable area segmentation and lane line segmentation
@@
Numerous methods have been proposed for individual tasks in autonomous driving, and many of them achieve outstanding results. For object detection tasks, two primary methodologies are prevalent. The first category is two-stage techniques, epitomized by Fast R-CNN [6]. While these methods prioritize detection accuracy, they often come at the expense of computational efficiency. The second category is one-stage approaches, epitomized by the You Only Look Once (YOLO) series [7, 8, 9, 10], ranging from version 1 to the latest, version 8. YOLOv8 is particularly noteworthy for its real-time detection capabilities, which have been widely adopted for various detection tasks. It primarily focuses on detection rather than segmentation. Although a segmentation head is introduced for segmentation tasks, the loss functions and evaluation metrics employed are mainly for the detection task. Additionally, YOLOv8 can only implement one task per model. Moreover, YOLOv8 is inherently limited in its ability to handle multiple tasks within the context of autonomous driving. Attempting to apply it under multi-task conditions necessitates the deployment of multiple models, leading to a substantial increase in both training and inference time.
@@
For the segmentation task, one of the field’s milestones is the advent of Fully Convolutional Networks (FCN) [11]. Additionally, U-Net [12] and SegNet [13] are common models in drivable area segmentation tasks. The drivable area typically covers a large portion of an image. Therefore, segmenting drivable areas requires information from high-level features. However, lane line segmentation differs from drivable area segmentation because of the distinct elongated and narrow features of lane lines in road images. Lane line segmentation needs low-level features for effective analysis. Recently, models such as PointLaneNet [14] and MFIALane [15] have gained popularity in the field of lane line segmentation. Although they achieve remarkable performance in segmentation tasks, the integration of these tasks with detection tasks into a single model poses challenges, primarily due to the differing resolution of features required. Segmentation operates at the pixel level, whereas object detection employs grid cells in one-stage methods and utilizes selective search in two-stage approaches. While their focus differs, both segmentation and detection tasks necessitate initial feature extraction from input images. Consequently, it is possible for them to share a common backbone. Compared to using separate models for each task, integrating three distinct necks and heads into a unified model with a shared backbone can significantly save computing resources and reduce inference time.
@@
Recently, several multi-task models have been proposed for ADS, such as YOLOP [16], multi-task learning model [17], Sparse U-PDP [18] and HybridNet [19]. All of these methods selected three tasks to construct the panoptic autonomous driving system: object detection, drivable area segmentation, and lane line segmentation, which are from the public Berkeley Deep Drive(BDD100K) dataset. Each of these methods has yielded outstanding results. Typically, their models are composed of two components: an encoder and a decoder. However, several challenges remain to be addressed. First, they have complex structures in their neck or head components, which can further impact inference times. For example, HybridNet [19] achieves outstanding performance in the detection task due to its use of an anchor-based head. This detection head performs better than the anchor-free head but tends to increase the inference time. Second, they often focus on specialized tasks when designing the loss function [16, 17] or neck structure [1]. However, this approach may compromise the model’s generality and necessitate considerable time for designing the structure and loss functions, as well as tuning their parameters. Therefore, developing a fast, robust, and universally applicable model is crucial.
@@
In this work, we propose A-YOLOM, an adaptive model specifically designed for multi-task. Notably, A-YOLOM can efficiently handle multi-task using a single model with a reasonable parameter overhead. This efficiency is due to the lightweight head we design for segmentation tasks, only built by a series of convolutional layers. Compared to other state-of-the-art multi-task methods, our model has fewer total parameters. Furthermore, by employing consistent loss functions across the same task types, we maintain a uniform approach to avoid customizing for specific tasks. Importantly, we introduce an adaptive module tailored for the segmentation neck. This is primarily achieved through a learnable parameter, trained to determine whether to concatenate features of different levels. This eliminates the need for distinct structural designs for various scene tasks. To summarize, the primary contributions of our study are as follows:
@@
  • We develop a lightweight model capable of integrating three tasks into a single model. This is particularly beneficial for multi-task that demand real-time processing, thereby enhancing the model’s deployability on edge devices.
@@
  • We design a novel Adaptive Concatenate Module specifically for the neck region of segmentation architectures. This module can adaptively concatenate features without manual design and achieve a similar or better performance than well-design, further enhancing the model’s generality.
@@
  • We design a lightweight, simple, and generic segmentation head. We adopt a unified loss function for the same type of task head, meaning we don’t need to custom design for specific tasks. It is only built by a series of convolutional layers.
@@
  • Extensive experiments are conducted based on publicly accessible autonomous driving datasets, which demonstrate that our model can outperform existing works, particularly in terms of inference time and visualization. Moreover, we further conduct experiments using real road datasets, which also demonstrate that our model significantly outperformed the state-of-the-art approaches.
@@ II Related Works
In this section, we review existing works on detection, segmentation, and multi-task models in autonomous driving tasks. We focus our discussion on methods based on deep learning.
@@ II-A Detection
Over the past decade, the swift advancements in computer vision have significantly bolstered the progress of autonomous driving. Specifically, autonomous driving tasks can be delineated into two primary parts: detection and segmentation. The detection task encompasses object detection, identifying entities such as vehicles, pedestrians, traffic signs, etc. The current object detection methods can be divided into two classifications [20]: two-stage and one-stage approaches.
Two-stage methods initiate with a Region Proposal Network (RPN) to generate Regions of Interest (RoI). Subsequently, the second phase employs a deep learning network to classify these RoIs. This latter stage also fine-tunes the bounding box dimensions and position, bolstering object localization accuracy. Noteworthy examples of two-stage object detection methods include Regions with CNN features (R-CNN) [21], Fast R-CNN [6], and Faster R-CNN [22].
In contrast, the one-stage detection approach offers an end-to-end strategy. It simultaneously predicts the bounding box and classifies the object in a single forward pass. It can be deployed on mobile devices for real-time operation and is easily implementable. The representatives of one-stage methods include You Only Look Once (YOLO) [7, 8, 9, 10] series, Single-Shot Multibox Detector (SSD) [23], and RetinaNet [24].
While one-stage methods typically lag behind two-stage methods in terms of detection performance, their real-time capabilities have made them increasingly popular in object detection tasks.
@@ II-B Segmentation
The segmentation task, including semantic and instance segmentation, differs from detection as it typically operates at the pixel level. Common segmentation tasks in autonomous driving include drivable area segmentation and lane line segmentation. However, these tasks exhibit notable differences. The drivable area typically covers a large portion of an image, but the number of regions is infrequent. As in Figure 1, only three drivable areas. Lane lines are the opposite: there are numerous in one image, yet each line is small, elongated, and narrow.
In 2015, the deep learning model Fully Convolutional Networks (FCN) [11] was introduced for semantic segmentation tasks. With its end-to-end training capability, it achieved a 20% improvement on the Pascal VOC 2012 dataset compared with traditional methods. However, its focus on local information leads to the loss of global details, thereby producing coarse segmentation results. To address this limitation, the encoder-decoder architecture model was proposed. Badrinarayanan et al. [13] proposed SegNet, which preserves the maximum pooling indices to ensure more accurate detail restoration during the upsampling process. However, the encoder-decoder architecture has its own shortcomings. High-resolution details are often lost during the encoding process, reducing fine-grained information. This is not ideal for tasks such as lane line segmentation. Pan et al. [25] proposed SCNN, which employs spatial convolutions to capture the continuity and structural information in images across both horizontal and vertical directions. Due to its ability to capture spatial correlations in images, it is well-suited for detecting slender objects, such as lane lines. However, such convolution operations are computationally intensive and time-consuming.
Given the constraints of edge devices in vehicles, it’s crucial to integrate multi-tasking into a single model. This not only conserves computational resources but also meets real-time performance requirements.
@@ II-C Multi-task model
Recently, multi-task models have gained considerable attention in the research community due to their high efficiency. This is particularly important for autonomous driving, which includes various sub-tasks and often operates under the constraints of limited computational resources on edge devices. MultiNet [26] introduced an efficient, unified deep architecture capable of simultaneously addressing classification, detection, and semantic segmentation. YOLOP [16] extended the YOLOv5 model by introducing two additional segmentation necks and heads to perform both segmentation and detection tasks simultaneously. They separately designed the loss function for each segmentation task, which compromises the generality of the model. Sparse U-PDP [18] proposed a unified decoding framework that integrates three tasks into a single model. Notably, they crafted a streamlined multi-task representation through “dynamic convolution kernels”, enhanced by a dynamic interaction module which adjusts feature sampling uniquely for each task. Due to its complex decoder, it is hard to meet real-time requirements. Shokhrukh et al. [17] introduced a memory-efficient end-to-end framework that utilizes combinations of customized loss functions, encompassing the weighted average sum of all tasks to enhance performance. The customization and specific weighting limited the model’s adaptability to new tasks. Additionally, their approach involved a sophisticated training paradigm, which could lead to overfitting and suboptimal real-world performance.
Therefore, the challenge of developing a generic and real-time multi-task model remains significant. In our work, we introduce an adaptive concatenation module and use the same loss function for tasks of the same type, further enhancing the model’s generality. To decrease parameter count, we employ a lightweight backbone and a simple yet effective segmentation head, ensuring the model meets real-time requirements under limited computing resources.
@@ III PROPOSED METHODOLOGY
In this section, we present the details of our model, including three main components: the backbone, neck, and head. Additionally, the loss function is also included. As shown in Fig 2. A-YOLOM model is a one-stage network with a simple encoder-decoder architecture. The encoder comprises both the backbone and neck, while the decoder consists of the head. It’s worth noting that we have three necks in total: one detection neck for object detection and two segmentation necks, one for drivable areas segmentation and the other for lane lines segmentation. Our model boasts an adaptive, simple and efficient structure, which not only broadens its application range but also ensures real-time inference.
@@ III-A Encoder
In our work, we integrate a shared backbone network and three neck networks into a single model for the three different tasks.
@@ III-A1 Backbone
The backbone comprises a series of convolutional layers to extract features from the input data. Due to the outstanding performance of YOLOv8 in detection tasks, our backbone network follows YOLOv8. Specifically, they refine CSP-Darknet53 [27], which serves as the backbone of YOLOv5. The key difference between this backbone and CSP-Darknet53 lies in the replacement of the c3 module with the c2f module. The c2f module combines high-level features with contextual information, further improving performance. There are also minor differences, such as using a 3x3 convolutional in place of the initial 6x6 convolutional, and the removal of the 10th and 14th convolutional layers. These enhancements make the backbone more effective than its predecessors in the YOLO series.
@@ III-A2 Neck
The neck is responsible for fusing the features extracted from the backbone. After the backbone network, we use a Spatial Pyramid Pooling Fusion (SPPF) module to increase the receptive field and reduce computational requirements compared to SPP. These features are subsequently directed to individual necks.
In our model, we utilize three necks: one for object detection task and two for segmentation tasks, which are drivable areas and lane lines. It’s worth noting that in Fig 2, as we have two segmentation tasks, the value of N in the segment neck on the left part is equal to 2. This means we have two same segment necks for different segmentation tasks. The sky-blue and brown lines, which originate from different neck bottom layers, lead to separate segment heads.
For the detection neck, we adopt the Path Aggregation Network (PAN) [28] structure, which includes top-down and bottom-up two Feature Pyramid Networks (FPN) [29]. This structure merges low-level details with high-level semantic features, enriching the overall feature representation. This is invaluable for object detection, where small objects hinge on low-level features, while larger entities gain from the broader context provided by high-level features. In detection tasks, objects of varying scales are commonly present. This diversity in object size is the primary reason we selected the PAN structure for the detection neck.
For the segmentation neck, we adopt the FPN structure, renowned for its effective handling of multi-scale objects. Additionally, we introduce an adaptive concatenation module between the neck and backbone. As illustrated in Algorithm 1, input includes two feature maps: one from the previous layer neck and the other from the identical resolution backbone. Central to this module is a learnable parameter, which aims to control the concatenation of backbone features, thereby enabling adaptive modifications to the model’s structure to accommodate different segmentation tasks. This module operates on each resolution level. This addition enhances the model’s generality, making it easily applicable across different segmentation tasks without customization.
@@
0:  List of tensors $x$ includes two feature maps, x[0] from neck, x[1] from identical resolution level backbone; channels list $ch$